OpenStudy (anonymous):

valuate the expression if a = 4 , b = 9. and c = 1. 3b(divided)c ----------- 1-a

OpenStudy (anonymous):

so all you need to do is substitute in your values 3b(divided)c = 3*9 div 1 something divided by 1 = the original something so 3*9 div 1 = 3*9 now 1-a, if a = 4 =1-4 so final equation is 3*9 ------------- 1-4 can you solve from here? @Mattfish21 ?